OpenStudy (bruno102):

A bag has 2 green marbles, 3 red marbles, and 3 yellow marbles. What is the probability that you pick a green marble, do not replace it, and pick a yellow marble?

OpenStudy (bruno102):

I keep getting 1/15 but the answer choices are 3/28, 1/28, 9/56 or 1/8

OpenStudy (anonymous):

its 2/8 * 3/7 = 6/56 = 3/28

OpenStudy (bruno102):

Ohh my bad, I was using the total number of marbles from a previous question i had. Thanks!

OpenStudy (anonymous):

:)